# Assumption-lean and Data-adaptive Post-Prediction Inference

### v3

@@ 2.2 Related work
Our setting is closely related to a recent method called prediction-powered (PP) inference (Angelopoulos et al., 2023a, ), which elicits valid inference with gold standard labels and arbitrary ML predictions. It corresponds to a special case of our POP-Inf estimator where $\omega=1$ (or a vector of all one’s). Both theoretical and numeric comparisons suggest that the PP estimator is less efficient than the proposed POP-Inf estimator.
@@
Our proposal of boosting an estimator with an unbiased estimator of zero has been a long-standing idea. One famous estimator is the augmented inverse probability weighting estimator in the literature of missing data (Robins et al.,, 1994) and causal inference (Robins,, 2000), in which a consistent inverse probability weighting estimator is augmented by a weighted residual term with mean zero, resulting in efficiency gain (Hahn,, 1998). In fact, it reaches the semiparametric efficiency bound (Bickel et al.,, 1993; Hahn,, 1998) when nuisance functions are correctly specified or approximated sufficiently fast. Although the pursuit of efficiency under correct specifications differs from our main focus, we establish the efficiency bound of our estimation problem (1) in Section 4 and illustrate the connection between the POP-Inf estimator and the efficient influence function in an example in Section 2.1.
@@
The idea of using unlabeled data fits into the broader context of improving efficiency with auxiliary data, which has been widely applied in survey estimation (Breidt and Opsomer,, 2017), missing data (Robins et al.,, 1994), measurement error models (Chen et al.,, 2005), causal inference with surrogate outcomes (Kallus and Mao,, 2020), and semi-supervised learning (Wang and Shen,, 2007; Chakrabortty and Cai,, 2018). The literature of semi-supervised learning has adopted a similar data structure with our problem in which labeled data are accompanied with unlabeled data (possibly with surrogate outcomes). Recently, Chakrabortty and Cai, (2018) have considered efficiency improvement on linear models under possible model misspecifications. After that, estimation methods have been proposed for mean estimation (Zhang et al.,, 2019), best linear predictor estimation (Azriel et al.,, 2022), and the general M-estimation problems (Song et al.,, 2023). Our work is distinct from this line of research in that we consider post-prediction inference with any “black-box” ML predictions. We aim at efficiency improvement with no assumptions on $\widehat{f}$ and for generally defined parameters.
3 POP-Inf
@@
In this section, we introduce an assumption-lean and data-adaptive post-prediction inference procedure that guarantees variance reductions for a general estimand defined in (1). We study asymptotic properties of the proposed POP-Inf estimator in Section 3.1 and provide an estimation algorithm in Section 3.2. Examples on regression models are provided in Section 3.3.
@@
Bearing in mind that $\widehat{f}$ could be incorrect, one of the most straightforward approaches to estimating ${\bm{\theta}}$ is simply ignoring unlabeled data and use labeled data only. This classic estimator $\widehat{\bm{\theta}}_{\textnormal{C}}$ solves for
$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})=0.$ (9)
It is always consistent, but can be of low efficiency when the labeled data is of a small size.
@@
In this work, we aim to provide an estimator that is consistent and more efficient than the classic estimator, regardless of the quality of the $\widehat{f}$. In addition to the classic approach that takes average on labeled samples, we motivate our method with an augmented mean zero term that is indexed by a vector $\bm{\omega}$,
@@
$\bm{\omega}\odot\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})\right\},$ (10)
where $\odot$ is the element-wise multiplication. We refer to the vector $\bm{\omega}$ as a weighting vector as it controls the contribution of labeled outcomes and the arbitrary prediction $\widehat{f}$.
In light of this, we propose our POP-Inf estimator that solves the equation,
$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+\bm{\omega}\odot\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})\right\}={\bf 0},$ (11)
where $\bm{\omega}$ can be replaced with an estimate $\widehat{\bm{\omega}}$ aiming at minimizing the variance. We shall elaborate an optimal value in Section 3.1 and an estimation procedure in Section 3.2. For any weighting vector $\bm{\omega}$, the resulting estimator is denoted as $\widehat{\bm{\theta}}_{\textnormal{POPInf}}(\bm{\omega})$.
3.1 Asymptotic properties
@@
We study the asymptotic properties of the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{POPInf}}(\bm{\omega})$ with any weight $\bm{\omega}$. To establish its consistency and asymptotic normality, we introduce the following regularity conditions. Condition (C1) is a regularity assumption on the parameter of interest. Conditions (C2)-(C3) guarantee consistency. Condition (C4) is needed for asymptotic normality. All conditions are reasonable and standard (Van der Vaart,, 2000).
@@
    There exists a function $\bm{\varphi}$ such that $\mathbb{E}[\sup_{{\bm{\theta}}\in\Theta}\bm{\varphi}^{2}]<\infty$ and $\mathbb{E}\lVert\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{1})-\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{2})\lVert_{2}^{2}\leq\bm{\varphi}\lVert{\bm{\theta}}_{1}-{\bm{\theta}}_{2}\rVert^{2}$ for any ${\bm{\theta}}_{1}$ and ${\bm{\theta}}_{2}$ in a neighborhood of ${\bm{\theta}}_{0}$. Furthermore, $\mathbb{E}\{\bm{\psi}(y,{\bf x};{\bm{\theta}})\}$ is differentiable at ${\bm{\theta}}_{0}$ with a nonsingular derivative matrix.
Theorem 1 (Asymptotic Normality).
@@
Under Conditions (C1)-(C3), assuming $\frac{n}{N}\rightarrow\rho$ as $n\to\infty$ and $N\rightarrow\infty$, then the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{POPInf}}(\bm{\omega})$ converges to ${\bm{\theta}}_{0}$ in probability. Assuming additionally Condition (C4), we have that as $n\to\infty$ and $N\to\infty$,
@@
$\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}({\bm{\omega}})-{\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bm{\Sigma}}(\bm{\omega})),$ (12)
@@
where ${\bm{\Sigma}}(\bm{\omega})={\bf A}^{-1}{\bf V}(\bm{\omega}){\bf A}^{-1}$, ${\bf A}=\mathbb{E}[\partial\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0})/\partial{\bm{\theta}}]$, ${\bf V}(\bm{\omega})={\bf M}_{1}+\bm{\omega}\bm{\omega}^{\rm T}\odot({\bf M}_{2}+\rho{\bf M}_{3})-2{\bf I}_{q}\bm{\omega}^{\rm T}\odot{\bf M}_{4}$, ${\bf M}_{1}=\textnormal{Var}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{2}=\textnormal{Var}_{l}[\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{3}=\textnormal{Var}_{u}[\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{4}=\textnormal{Cov}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0}),\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$. Proof is contained in Appendix B.1.
Theorem 1 suggests that the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{POPInf}}({\bm{\omega}})$ is asymptotically normal with any weight vector $\bm{\omega}$. With the pursuit of minimizing the asymptotic variance, we define the optimal weight $\bm{\omega}_{0}=(\omega_{01},\cdots,\omega_{0q})^{\rm T}$ such that each element $\omega_{0j}$ minimizes the variance of the corresponding element in the parameter,
$\omega_{0j}=\arg\min\limits_{\omega_{j}\in\mathbb{R}}{\bm{\Sigma}}_{j}{(\bm{\omega})}\text{ for }j=1,\cdots,q,$ (13)
where ${\bm{\Sigma}}_{j}(\bm{\omega})$ is the $j$-th diagonal element of the covariance matrix ${\bm{\Sigma}}{(\bm{\omega})}$. A data adaptive estimation procedure for $\bm{\omega}$ is provided in Section 3.2.
Remark 3 (Variance Comparison with the Classic Approach).
The asymptotic variance in Theorem 1 implies a direct comparison with the classic approach with labeled data only. Let $v_{j}(\omega_{j})$ be the $j$-th diagonal element of matrix ${\bf V}(\bm{\omega})$, then
$v_{j}(\omega_{j})=m_{1j}+(m_{2j}+\rho m_{3j})\omega_{j}^{2}-2m_{4j}\omega_{j},$ (14)
where $m_{1j}$, $m_{2j}$, $m_{3j}$ and $m_{4j}$ are the $j$-th diagonal element of matrices ${\bf M}_{1}$, ${\bf M}_{2}$, ${\bf M}_{3}$ and ${\bf M}_{4}$, respectively. The classic approach with $\omega_{j}=0$ corresponds to
$v_{j}(0)=m_{1j}.$ (15)
We compare it with the minimum of $v_{j}(\omega_{j})$ at two cases of $\rho$, the limit of $n/N$.
  • When $\rho\in[0,\infty)$, the minimum of $v_{j}(\omega_{j})$ is no greater than $v_{j}(0)$,
    $\min_{\omega_{j}}v_{j}(\omega_{j})=v_{j}(m_{4j}/(m_{2j}+\rho m_{3j}))$ $=m_{1j}-\dfrac{m_{4j}^{2}}{m_{2j}+\rho m_{3j}}\leq m_{1j},$ (16)
    suggesting the pursuit of a smaller variance. Note that $\rho\in[0,\infty)$ contains the special case $\rho=0$ which is common in the literature of semi-supervised learning.
  • When $\rho=\infty$, that is when $N=0$, we have $\min_{\omega_{j}}v_{j}(\omega_{j})=v_{j}(0)$. This suits our intuition that when all data are labeled, the optimal choice is asymptotically as efficient as the classic approach.
Remark 4 (Data-adaptive Variance Reduction).
In (14), we notice that the last term is $-2m_{4j}\omega_{j}$. Recall that $m_{4j}$ is the $j$-th diagonal entry of the covariance matrix between true outcomes and predicted outcomes on the labeled data, it reflects the prediction accuracy: higher $m_{4j}$ implies higher prediction accuracy. If $\omega_{j}$ is fixed and nonzero, then $-2m_{4j}\omega_{j}$ becomes smaller as $m_{4j}$ increases and vice versa, indicating that a better prediction algorithm yields a smaller variance. While this observation should hold in general, our proposed POP-Inf method allows a data-adaptive choice of $\omega_{j}$ so that it amplifies variance reduction when the prediction performs well, and avoids variance inflation otherwise.
Remark 5 (Element-wise Variance Reduction).
The optimal choice weight (13) is defined for every element of the parameter so that each component is targeted to have a reduced variance. Under cases where only a subset of the parameter is of inferential interest, it’s possible to only assign nontrivial weights to corresponding elements. For example in a regression model, if one only aims to make inference on the coefficient of a particular covariate, then the other $\omega_{j}$’s can be set to zero.
The asymptotic normality established in Theorem 1 enables variance estimation and statistical inference. Corollary 1 illustrates the construction of a confidence interval, where a consistent estimate $\widehat{\bm{\Sigma}}{(\bm{\omega})}$ of the covariance matrix can be easily obtained by the sample counterpart of ${\bm{\Sigma}}{(\bm{\omega})}$.
Corollary 1 (Confidence Interval).
Suppose $\widehat{\bm{\Sigma}}{(\bm{\omega})}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\Sigma}}{(\bm{\omega})}$. Given confidence level $1-\alpha\in(0,1)$, construct an interval for the $j$-th element of the parameter, $\theta_{0,j}$, as
${\cal C}_{\alpha,j}=\left\{\theta:\theta\in\left(\widehat{\theta}_{\textnormal{POP-Inf}}({\bm{\omega}})_{j}-z_{\alpha/2}\{\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}/n\}^{-1/2},\widehat{\theta}_{\textnormal{POP-Inf}}({\bm{\omega}})_{j}+z_{\alpha/2}\{\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}/n\}^{-1/2}\right)\right\},$ (17)
where $\widehat{\theta}_{j}$ is the $j$-th element of $\widehat{\bm{\theta}}_{\textnormal{POPInf}}$ and $z_{\alpha/2}$ is the upper $\alpha/2$ quantile of the standard normal distribution. Then
$\lim_{N\in\infty}\mathbb{P}(\theta_{0,j}\in{\cal C}_{\alpha,j})\geq 1-\alpha,\ j=1\cdots,q.$ (18)
Hence, ${\cal C}_{\alpha,j}$ is a confidence interval for $\theta_{0,j}$ at level $1-\alpha$.
3.2 Estimation procedure
In practice, we propose to estimate the weight with $\widehat{\bm{\omega}}$ that minimizes each element of the sample counterpart of the covariance matrix. That is,
$\widehat{\omega}_{j}=\arg\min\limits_{\omega_{j}\in{\bf R}}\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}\text{ for }j=1,\cdots,q,$ (19)
where $\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}$ is the $j$-th diagonal element of the sample analog of ${\bm{\Sigma}}{(\bm{\omega})}$.
The following Algorithm 1 details an adaptive estimation procedure that updates the estimates for both the weight and the parameter of interest at each iteration. In the end, after obtaining the estimator and the corresponding weight, one can construct a confidence interval as stated in Corollary 1. Note that this algorithm is applicable to any parameter defined in (1). It can be simplified under certain scenarios. As illustrated in Section 2.1, no iterations are needed to estimate the outcome. In that example, one can consistently estimate the weight directly from data and build an estimation equation for estimating the outcome mean.
Algorithm 1 POP-Inf Estimation
1:Data ${\cal L}\cup{\cal U}$, predictor $\widehat{f}$, $\bm{\omega}^{0}={\bf 0}$, error rate $\alpha\in(0,1)$, a step size $\eta$ for gradient descent, tolerance $\delta$, and the maximum of iterations $K$.

### v4

@@ 2.2 Related work
Our setting is closely related to a recent method called prediction-powered inference (PPI) (Angelopoulos et al., 2023a, ), which ensures valid inference with gold-standard labels and arbitrary ML predictions. It corresponds to a special case of our PSPA estimator where $\omega=1$ (or a vector of all ones). Both theoretical and numeric comparisons suggest that the PPI estimator is less efficient than the proposed PSPA estimator. Our approach is also closely related to a concurrent method called PPI++ (Angelopoulos et al., 2023b, ), which also aims to enhance the efficiency of PPI. PSPA is guaranteed to be more efficient than PPI++. Unlike PSPA, PPI++ cannot guarantee element-wise variance reduction compared with the classical method in multi-dimensional estimation problems. Additionally, PPI++ cannot be applied to statistical inference with ML-predicted covariates, while PSPA is capable of handling this type of applications. We present a comprehensive comparison of PSPA and PPI++ in Section 8.
@@
Our proposal of boosting a consistent estimator with a consistent estimator of zero has been a long-standing idea. One famous estimator is the augmented inverse probability weighting estimator in the literature of missing data (Robins et al.,, 1994) and causal inference (Robins,, 2000), in which a consistent inverse probability weighting estimator is augmented by a weighted residual term with mean zero, resulting in efficiency gain (Hahn,, 1998). In fact, it reaches the semiparametric efficiency bound (Bickel et al.,, 1993; Hahn,, 1998) when nuisance functions are correctly specified or approximated sufficiently fast. Although the pursuit of efficiency under correct specifications differs from our main focus, we establish the efficiency bound of our estimation problem (1) in Section 4 and illustrate the connection between the PSPA estimator and the efficient influence function.
@@
The idea of using unlabeled data fits into the broader context of improving efficiency with auxiliary data, which has been widely applied in survey estimation (Breidt and Opsomer,, 2017), missing data (Robins et al.,, 1994), measurement error models (Chen et al.,, 2005), causal inference with surrogate outcomes (Kallus and Mao,, 2020), and semi-supervised learning (Wang and Shen,, 2007; Chakrabortty and Cai,, 2018). The literature on semi-supervised learning has adopted a similar data structure to our problem in which labeled data are accompanied by unlabeled data (possibly with surrogate outcomes). Recently, Chakrabortty and Cai, (2018) have considered efficiency improvement in linear models under possible model misspecifications. After that, estimation methods have been proposed for mean estimation (Zhang et al.,, 2019), best linear predictor estimation (Azriel et al.,, 2022), and general M-estimation problems (Song et al.,, 2023). Our work is distinct from this line of research in that we consider post-prediction inference with any “black-box” ML predictions for efficiency improvement.
3 PSPA for assumption-lean and data-adaptive post-prediction inference
@@
In this section, we introduce our method PSPA for assumption-lean and data-adaptive post-prediction inference. For estimand defined by (1) and any ML prediction, it guarantees the validity of inference results and element-wise variance reduction compared with the classical method that uses labeled data alone. We present our estimator and algorithm in Section 3.1 and establish the theoretical guarantees for our method in Section 3.2. Examples of applying our method for regression tasks are provided in Section 3.3.
3.1 PSPA estimator
@@
Bearing in mind that $\widehat{f}$ could be incorrect, one typical approach to estimating ${\bm{\theta}}$ is to ignore the unlabeled data and use only the labeled data. This classical estimator $\widehat{\bm{\theta}}_{\textnormal{C}}$ solves for
$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})={\bf 0}.$ (9)
It is always consistent, but may have low efficiency because it ignores the unlabeled data.
@@
In this work, we aim to provide an estimator that is consistent and more efficient than the classical estimator, regardless of the quality of the $\widehat{f}$. Similar to the mean estimation example, we add an augmentation term, which has a mean of zero and is indexed by a vector ${\bm{\omega}}=[\omega_{1},\cdots,\omega_{q}]^{\rm T}$, to the estimation equation for the classical approach:
${\rm diag}({\bm{\omega}})\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f},{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f},{\bf x}_{i};{\bm{\theta}})\right\},$ (10)
where ${\rm diag}({\bm{\omega}})$ is a diagonal matrix with diagonal elements $\omega_{1}$, $\cdots$, $\omega_{q}$. We refer to the vector ${\bm{\omega}}$ as a weighting vector as it controls the contribution of labeled outcomes and the arbitrary prediction $\widehat{f}$ in a data-adaptive way. Together, we propose our PSPA estimator $\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}({{\bm{\omega}}})$ that solves the equation,
@@
$\Psi_{\textnormal{{PSPA}}}^{\bm{\omega}}({\bm{\theta}}):=\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+{\rm diag}({\bm{\omega}})\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f},{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f},{\bf x}_{i};{\bm{\theta}})\right\}={\bf 0},$ (11)
@@
for a given weighting vector ${\bm{\omega}}$. This equation corresponds to a class of estimators $\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}({{\bm{\omega}}})$ for different fixed values of ${\bm{\omega}}$. We next establish the consistency and asymptotic normality of these estimators with the following regularity conditions. Condition (C1) is a regularity assumption on the parameter of interest. Conditions (C2)-(C4) guarantee consistency. Condition (C5) is needed for asymptotic normality. All conditions are reasonable and standard (Van der Vaart,, 2000).
@@
    Labeled and unlabeled data are i.i.d drawn from the population of interest.
  • (C5)
@@
    There exists a function $\bm{\varphi}$ such that $\mathbb{E}[\sup_{{\bm{\theta}}\in\Theta}\bm{\varphi}^{2}]<\infty$ and $\mathbb{E}\lVert\bm{\psi}(Y,{\bf X};{\bm{\theta}}^{\prime})-\bm{\psi}(Y,{\bf X};{\bm{\theta}}^{\prime\prime})\lVert_{2}^{2}\leq\bm{\varphi}\lVert{\bm{\theta}}^{\prime}-{\bm{\theta}}^{\prime\prime}\rVert^{2}$ for any ${\bm{\theta}}^{\prime}$ and ${\bm{\theta}}^{\prime\prime}$ in a neighborhood of ${\bm{\theta}}$. Furthermore, $\mathbb{E}\{\bm{\psi}(y,{\bf x};{\bm{\theta}})\}$ is differentiable at ${\bm{\theta}}$ with a nonsingular derivative matrix.
Theorem 1.
@@
Under Conditions (C1)-(C4), assuming $\frac{n}{N}\rightarrow\rho$ as $n\to\infty$ and $N\rightarrow\infty$, then the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}({{\bm{\omega}}})$ converges to ${\bm{\theta}}$ in probability. Assuming additionally Condition (C5), we have
@@
$\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}({{\bm{\omega}}})-{\bm{\theta}})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bm{\Sigma}}({\bm{\omega}})),$ (12)
@@
where ${\bm{\Sigma}}({\bm{\omega}})={\bf A}^{-1}{\bf V}({\bm{\omega}}){\bf A}^{-1}$, ${\bf A}=\mathbb{E}[\partial\bm{\psi}(Y,{\bf X};{\bm{\theta}})/\partial{\bm{\theta}}]$, ${\bf V}({\bm{\omega}})={\bf M}_{1}+{\rm diag}({\bm{\omega}})({\bf M}_{2}+\rho{\bf M}_{3}){\rm diag}({\bm{\omega}})-2{\rm diag}({\bm{\omega}}){\bf M}_{4}$, ${\bf M}_{1}=\operatorname{Var}[\bm{\psi}(Y,{\bf X};{\bm{\theta}})]$, ${\bf M}_{2}=\operatorname{Var}[\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}})]$, ${\bf M}_{3}=\operatorname{Var}[\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$,
${\bf M}_{4}=\operatorname{Cov}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}),\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}})]$.
The proof is contained in Appendix B.1. Theorem 1 suggests that the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{{PSPA}}}({{\bm{\omega}}})$ is asymptotically normal with a fixed weighting vector ${\bm{\omega}}$. In practice, this weighting vector ${\bm{\omega}}$ should be estimated from data with the pursuit of efficiency. We provide an estimation procedure aimed at element-wise variance reduction. This is made possible by a key observation that the $j$-th diagonal element of the asymptotic covariance matrix ${\bm{\Sigma}}({\bm{\omega}})$ is a quadratic function of $\omega_{j}$ and does not depend on other components of ${\bm{\omega}}$:
${\bm{\Sigma}}_{jj}({\bm{\omega}})=\omega_{j}^{2}[{\bf A}^{-1}({\bf M}_{2}+\rho{\bf M}_{3}){\bf A}^{-1}]_{jj}-2\omega_{j}[{\bf A}^{-1}{\bf M}_{4}{\bf A}^{-1}]_{jj}+[{\bf A}^{-1}{\bf M}_{1}{\bf A}^{-1}]_{jj}$ (13)
where $[{\bf M}]_{jj}$ represents the $j$-th diagonal element of matrix ${\bf M}$. We define the optimal weighting vector ${\bm{\omega}}^{\textnormal{opt}}=[\omega_{1}^{\textnormal{opt}},\dots,\omega_{q}^{\textnormal{opt}}]^{\rm T}$ such that each coordinate of the weighting vector minimizes the asymptotic variance for the corresponding coordinate:
$\omega_{j}^{\textnormal{opt}}$ $=\arg\min_{\omega_{j}}{\bm{\Sigma}}_{jj}(\omega_{j})=\dfrac{[{\bf A}^{-1}{\bf M}_{4}{\bf A}^{-1}]_{jj}}{[{\bf A}^{-1}({\bf M}_{2}+\rho{\bf M}_{3}){\bf A}^{-1}]_{jj}}\text{ for all }j\in 1,\dots,q,$ (14)
The resulting asymptotic variance for $j$-th coordinates is
${\bm{\Sigma}}_{jj}({\bm{\omega}}^{\textnormal{opt}})$ $=[{\bf A}^{-1}{\bf M}_{1}{\bf A}^{-1}]_{jj}-\dfrac{[{\bf A}^{-1}{\bf M}_{4}{\bf A}^{-1}]_{jj}}{[{\bf A}^{-1}({\bf M}_{2}+\rho{\bf M}_{3}){\bf A}^{-1}]_{jj}}.$ (15)
Such construction of ${\bm{\omega}}^{\textnormal{opt}}$ guarantees the optimality of the element-wise asymptotic variance across the class of estimators with any weighting vector ${\bm{\omega}}$. Next, we present our algorithm for estimation and inference.
Algorithm 1 PSPA estimation with ML-predicted labels
1:Data ${\cal L}\cup{\cal U}$, pre-trained ML model $\widehat{f}$, error rate $\alpha\in(0,1)$.